OpenStudy (anonymous):

My professor converted 8pi/3 rdn to -2pi - 2pi/3. Can anyone explain how he got that? Is there a certain formula for it? The problem I'm on now is -31pi/4

OpenStudy (anonymous):

\[-\frac{8\pi}{3} = -\frac{6\pi+ 2\pi}{3}= -\frac{6\pi}{3}-\frac{ 2\pi}{3}=-2\pi-\frac{ 2\pi}{3}\]

OpenStudy (anonymous):

Thanks! the only question I have still is do I want to limit the pi in the numerator to 2pi and just subtract all remaining pi to the left?

OpenStudy (anonymous):

I didn't understand your question :) but what I think you should do here is find the largest number that is smaller than 31 but is still divisible by 4, then split it to have only pi, and a remainder (something*pi/4). show me your answer and I'll give you my opinion

OpenStudy (anonymous):

Okay, I ended up with converting it to 28pi+3pi/4 = 28pi/4 - 3pi/4 = 7pi -3pi/4. I think thats the final answer (sorry for the bad formatting, still not used to using symbols in this chat)